OpenStudy (anonymous):

Which is an equation of the line passing through points at (1, 6) and (-2, 1)? y = 2x + 5 y = 5/3x + 13/3 y = x + 5 y = -1/2x +13/2

OpenStudy (anonymous):

Try getting the slope of the line first, using the formula: \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

The slope is -5/3

OpenStudy (anonymous):

i mean -5/-3

OpenStudy (anonymous):

Great, that's what I got as well. Those minus signs will cancel each other out too, and you'll have \(\frac{5}{3}\) as the slope. The next step is to plug some numbers into the equation for a line, which is: \(y-y_1 = m(x-x_1)\)

OpenStudy (anonymous):

so the answer is y = 5/3x + 13/3

OpenStudy (anonymous):

That's what I got, and it matches up to option B above. Congrats on getting it right!